# Elastic scattering at $\sqrt{s}=6$ GeV up to $\sqrt{s}=13$ TeV

### v1

@@ Abstract
In the framework of the Regge-eikonal model of hadron interaction, based on the analyticity of the scattering amplitude with taking into account the hadron structure, the simultaneous analysis is carried out of the 90 sets of data. These sets include the data obtained at low energies ($\sqrt{s}>3.6$ GeV and at high energies at FNAL, ISR, $S\bar{p}pS$, TEVATRON and LHC with 4326 experimental points, including the polarization data of analysing power. The energy and momentum transfer dependence of separate sets of data is analyzed on the basis of the eikonalized Born amplitude with taking into account two additional anomalous terms. Different origins of the nonlinear behavior of the slope of the scattering amplitude are compared. No contribution of hard-Pomeron in the elastic hadron scattering is shown. However, the importance of Odderon’s contribution is presented. The form and energy dependence of different terms of the cross-even and cross-odd parts of the elastic nucleon-nucleon scattering amplitude is determined in the framework of the High Energy Generalize Structure (HEGS) model. In the framework of the HEGS model, using the electromagnetic and gravitomagnetic form factors, the differential cross sections in the Coulomb Nuclear Interference (CNI) region and at large momentum transfer are described well in a wide energy region simultaneously. It is shown that the cross-even part includes the soft pomeron growing like $\ln^{2}(s)$ and an additional term with a large slope and with and with energy dependence $\ln^{2}(s)$. The cross-odd part includes the maximal odderon term and an additional oscillation term with $\ln(s)$. It is shown that both additional terms are proportional to charge distributions but the maximal odderon term is proportional to matter distributions. Also, a good description of proton-neutron differential scattering with 526 experimental points it is obtained on the basis of the amplitudes taken from the analysis of $pp$ and $p\bar{p}$ scattering. A good enough description of the polarization data was also obtained.
@@ I Introduction
One of the most important tasks of modern physics is the research into the basic properties of hadron interactions. Many models predict that soft hadron interactions will enter a new regime at the LHC: given the huge energy, as the S-matrix reaches the unitarity limit. The dynamics of strong interactions finds its most complete representation in elastic scattering. It is just this process that allows the verification of the results obtained from the main principles of quantum field theory: the concept of the scattering amplitude as a unified analytic function of its kinematic variables connecting different reaction channels was introduced in the dispersion theory by N.N. Bogoliubov[1]. Now many questions of hadron interactions are connected with modern problems of astrophysics such as unitarity and the optical theorem [2], and problems of baryon-antibaryon symmetry and CP-invariance violation [3] The main domain of elastic scattering is small angles. Only in this region of interactions we can measure the basic properties that define the hadron structure. Their values are connected, on the one hand, with the large-scale structure of hadrons and, on the other hand, with the first principles that lead to theorems on the behavior of scattering amplitudes at asymptotic energies [4, 7].
@@
The research of the structure of the elastic hadron scattering amplitude at superhigh energies and small momentum transfer - $t$ can give a connection between the experimental knowledge and the basic asymptotic theorems based on first principles [5, 6, 7]. It gives information about the hadron interaction at large distances where the perturbative QCD does not work [8, 9] and a new theory as, for example, instanton or string theories must be developed.
@@
The structure of hadrons reflected in generalized parton distributions is now one of the most interesting questions of the physics of strong interactions (see, for example [12, 15]. It is tightly connected with the spin physics of the hadron [16]. Some modern accelerator experiments have developed extensive programs for deep studies of different issues related to this problem. For example, the Jefferson Laboratory/Electron-Ion Collider (EIC) team to extract generalized parton distributions (GPDs) [17]. the same problem is posed for future experiments at SPD of NICA-JINR [18].
@@
Modern studies of elastic scattering of high energy protons lead to several unexpected results reviewed, e.g., in [9, 10]. Spin amplitudes of the elastic $NN$ scattering constitute a spin picture of the nucleon. Without knowing of the spin $NN$-amplitudes, it is impossible to understand the spin observable of nucleon scattering off nuclei. In the modern picture, the structure of hadrons is determined by Generalized Distribution functions (GPDs), which include the corresponding parton distributions (PDFs). The sum rules [20, 21] allows one to obtain the elastic form factor (electromagnetic and gravitomagnetic) through the first and second integration moments of GPDs. It leads to remarkable properties of GPDs, some corresponding to inelastic and elastic scattering of hadrons. Now some different models examining the nonperturbative instanton contribution lead to sufficiently large spin effects at superhigh energies [57], [56]. The research of such spin effects will be a crucial stone for different models and will help us to understand the interaction and structure of particles, especially at large distances. There are large programs of researching spin effects at different accelerators. Especially, we should like to note the programs at NICA, where the polarization of both the collider beams will be constructed. So it is very important to obtain reliable predictions for spin asymmetries at these energies. In this paper, we extend the model predictions to spin asymmetries in the NICA energy domain.
@@
There are two sets of data - at small momentum transfer [29] and at middle and large momentum transfer [30]. They overlap in some region of the momentum transfer, which supplies practically the same normalization of both sets of differential cross sections of elastic proton-proton scattering. Recently, the first set of data has created a wide discussion of the determination of the total cross section and the value of $\rho(t=0)$.
@@
There is a very important characteristic of the elastic scattering amplitude such as the ratio of the real part to imaginary part of the scattering amplitude - $\rho(s,t)$. It is tightly connected with the integral and differential dispersion relations. Of course, especially after different results obtained by the UA4 and UA4/2 Collaborations, physicists understand that $\rho(s,t=0)$ is not a simple experimental value but heavily dependent on theoretical assumptions about the momentum depends of the elastic scattering amplitude. Our analysis of both experimental data obtained by the UA4 and UA4/2 Collaborations shows a small difference value of $\rho(s,t=0)$ obtained in both the experiments if the nonlinear behaviour of the elastic scattering amplitude is taken into account [73]. Hence, this is not an experimental problem but a theoretical one [58].
@@
For extraction of the sizes of $\sigma_{tot}$ and $\rho(t=0)$ the Coulomb hadron region of momentum transfer is used (for example [29]). However, the form of the scattering amplitude assumed for small $t$ and satisfying the existing experimental data at small momentum transfer, can essentially be different from experimental data at large $t$. One should take into account the analysis of the differential cross section at 13 TeV where the diffraction minimum impacts the form of $d\sigma/dt$ already at $t=-0.35$ GeV2. The analysis of new effects, discovered on the basis of the experimental data at 13 TeV [58, 75, 76] and associated with the specific properties of the hadron potential at large distances was carried out with taking account all sets of experimental data on elastic $pp$-scattering obtained by the TOTEM and ATLAS Collaborations in a wide momentum transfer region and gave a quantitative description of all examined experimental data with minimum fitting parameters.
@@
The non-linear behavior of the slope of the differential cross sections at small momentum transfer, which was announced by the TOTEM Collaboration in the proton-proton elastic scattering at $8$ TeV, shows that the complex (complicated) form of strong interactions searched out at low energies remains at superhigh energies too. This means that the strong interaction is not simplified at superhigh energies but includes many different parts of hadron interactions.
@@
Using the existing model of nucleon elastic scattering at high energies $\sqrt{s}>3.6$ GeV - 14 TeV [38, 39], which involves minimum of free parameters, we are going to develop its extended version aimed to describe all available data on cross sections and spin-correlation parameters at lower energies down to the SPD NICA region. The model will be based on the usage of known information on GPDs in the nucleon, electro-magnetic and gravitomagnetic form factors of the nucleon taking into account analyticity and unitarity requirements and providing compatibility with the high energy limit, where the pomeron exchange dominates.
@@ II.1 Electromagnetic scattering
The question of the $t$ dependence of the elastic peak is one. One can be related to optics. At high energy, elastic scattering at small angles has some similarity with light scattering in the Fresnel region. The amplitude for the electric field propagating in the $z$ direction is given by the Rayleigh-Sommerfeld equation
@@
where $r=\sqrt{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}+z^{2}}$,and $\cos\theta=\frac{z}{r}$. The integral can be performed analytically only for the simplest geometrical cases, and one of the general forms of expansion over $r$ near the $z$ axis is
$r=$ $z\sqrt{1+\frac{\rho^{2}}{z^{2}}}=z\left[1+\frac{\rho^{2}}{2z^{2}}-\frac{1}{8}\left(\frac{\rho^{2}}{z^{2}}\right)^{2}+\cdots\right]$ (2)
$=$ $z+\frac{\rho^{2}}{2z}-\frac{\rho^{4}}{8z^{3}}+\cdots$
with $\rho^{2}=(x-x^{\prime})^{2}+(y-y^{\prime})^{2}$. If we consider only the first term and the case of a sphere of radius $R$, we obtain the $t$ dependence of the amplitude [33]
${\cal A}(t)\propto(1+\cos\theta)\frac{J_{1}(x\sin\theta)}{\sin\theta}$
with $x=kR$. Hence even in the simplest case the elastic peak cannot be described by a simple exponential. Note that this form of the scattering amplitude is used in diffractive hadron scattering.
@@ II.2 Hadron scattering
In reference [32], the TOTEM Collaboration has announced the observation of a the non-exponential behavior of the differential elastic cross sections at $8$ TeV and small momentum transfer $|t|$.
Of course, the form of the elastic peak depends also on the structure of particles and the dynamics of the interaction. These two features can be parameterised by the profile function $\rho(\vec{b})$ (in the space of impact parameter, $\vec{b}$) or by the elastic form factor (in momentum space). For example, in [34] different forms of the profile function determined by the density distributions taking part in the interaction were considered:
$\mathrm{circle\ of\ radius\ }b_{0}=4\mathrm{\ GeV^{-1}}$
@@
$\mathrm{hollow\ disk\ near\ }b_{0}=2.8\mathrm{\ GeV^{-1}}$
@@
blackdiskofradius $\mathrm{black\ disk\ of\ radius\ 6}\mathrm{\ GeV^{-1}}$
@@
Figure 1: Possible $t$ dependences of the scattering amplitude ${\cal A}(t)$, rescaled to 1 at $t=0$. (solid line - for Eq. (6); long-dashed line - for Eq. (3); short-dashed line - for Eq. (5); dotted line - for Eq. (4, and dashed-dotted line - for Eq. (7).
The $t$ dependence of these amplitudes is shown in Fig.1. Clearly, the simple exponential form (6) is a special separate case. We can see, except for the last case (eq.7), that if we take a small interval of $t$ (for example, $0.150.01<|t|<$ 0.15 GeV2) the curves give practically the same result. However, they give the essentially different results on some more values of $t$ or at $t\rightarrow 0$. Hence, to obtain the true form of the scattering amplitude, it is necessary to take experimental data in a wider interval of $t$ and tending to the limit $t\rightarrow 0$.
This simplest approximation of the differential cross section by one exponent with a constant slope can lead to artificial new effects. For example, in the experiment at the Protvino accelerator on the elastic proton-proton scattering there were found ”oscillations” at small momentum transfer. However, in [35] it was shown that such ”oscillations” can be appear when the differential cross section is described by two exponentials with some different slopes, by the one exponential with the constant slope.
@@ II.3 Non-linear slope
The complicated $t$ dependence of the slope can have many origins. First of all, it comes from the unitarization procedure of the Born term of the elastic scattering amplitude. In many purely phenomenological analyses is represented by the $Ct^{2}$ term in the slope that mimics the unitarization procedure. The hadron structure is reflected also not only at large momentum transfer but also at small $t$. This especially it is concerns the meson clouds whose interaction in many models is added to the central part of the hadrons (for example, in Pumplin model [36] and the Dubna Dynamical (DD) model [37])) which leads to the $\sqrt{t_{0}+t}$ dependence of the scattering amplitude. Such complicated hadron structure can be reflected in the presence of the two form factors - electromagnetic and mater form factor of the hadrons (for example, the HEGS model [38, 39]).
@@
used in many phenomenological descriptions of the elastic differential cross sections to explain the ”break” in the differential cross sections. Note that an additional term in the slope like $\sqrt{t_{0}-t}$ was early obtained whose first approximation can be related to absorbtion corrections that can produce a set of canceling Regge cuts [42, 43] and leads to the Schwarz type trajectories [44] $\alpha(t)=1+\gamma t^{1/2}$. A more complicated form was obtained in [45] $\alpha_{\pm}(t)=1\pm\gamma t^{1/2}+2\rho(1/2\ \gamma^{2}t)^{3/2}(-ln(t))^{1/2}$. The appearence of a complex trajectory greatly complicates the picture and requires additional research. Hence in [46], based on the works [54, 55], it was proposed to use the simplest form $\alpha(t)=1.041-0.15\sqrt{t_{0}-t}$. However, as we show such behavior has a really pure phenomenological basis and can be replaced by either a simpler a form (for example eq.(11) or more complicated form used in the HEGS-model [39].
Many models based on the famous works [47, 48] researched the non-linear behavior of the scattering amplitude.
Based on the works in [49] it was obtained
@@
with $q^{2}=-t$ Note that we have removed the misprint in this equation, as made in [51]. They obtained the limits of the representation in the brackets at $q^{2}>>4\mu^{2}$; hence, the slope grows in order $q^{2}$ with small logarithmic suppression. At small $t$ ($q^{2}\ll 4\mu^{2}$ they predicted that the representation in the brackets goes to $ln(m^{2}/\mu^{2})-8/3$. Note that the authors aimed to explain the deviation of the slope from the constant at non-small momentum transfer (in the domain $-t=0.4$ GeV2. However, in this domain the impact of the diffraction minimum is already felt. Hence this domain of $t$ is usually described by an additional term like $ct^{2}$, which were proposed early (for example [53]).
Figure 2: The $t$ dependence of the different parts of the terms of eq.(13) (solid line - the sum all three terms situated into brackets; the dots-dashed line the first term is multiplied by ($-1$); the long-dashed line - the second term, the short-dashed line - the sum of the two first terms).
Figure 3: The $t$ dependence of the additional terms to $\alpha^{\prime}t$ (solid line - eq.(13); long-dashed line - the eq.(8); short-dashed line - the eq.(8) without extraction $2\mu$ and multiplied on $0.1$.
@@
In [51], using the calculations [49] and removing the misprint, the authors proposed for $h_{1}$, with taking into account of the meson form factor with $\Lambda_{\pi}^{2}=m_{\rho}^{2}$ GeV2 ), the following equation:
$h_{1}(q^{2})=\frac{4}{\tau}f^{2}_{\pi}(t)[2\tau-(1+\tau)^{3/2}ln\frac{\sqrt{1+\tau}+1}{\sqrt{1+\tau}-1}+ln\frac{m^{2}}{\mu^{2}}],$ (13)
@@
Really for the slope, which is multiplied by $q^{2}$, there are two divergence terms with different signs plus a small constant term. The divergence terms cancels each other and the rest have a slow $t$ dependence (see Fig.2). The cancelation of the two terms as $t\rightarrow 0$ is not full and the rest have also some indefiniteness and strongly depend on $t$. In Fig.2, the $t$ dependence of the different parts of the term of eq.(13) in the brackets are shown. The cancelation of the two diverse terms leads to the negative contribution to the standard constant slope. However, the sum of all three terms give a positive additional contribution with small $t$ dependence.
In Fig. 3, we show the $t$ dependence of a different form of the slopes with the full kinematic coefficient. One can see very different $t$ dependence , especially in the form $\sqrt{4\mu^{2}+q^{2}}$ which was proposed in the work [46] without the extraction of $2\mu$.
@@
Some models of hadron interactions at high energy suppose that the slope has a slow $t$ dependence. For example, the Dubna dynamical model (DDm) [37], which takes into account the contribution to hadron interactions from the meson cloud of the nucleon and uses the standard eikonal form of the unitarization, leads to the scattering amplitude in the form
$T(s,t)=-is\sum_{n=1}^{\infty}\frac{\mu}{(n^{2}\mu^{2}-t)^{3/2}}(1-b\sqrt{n^{2}\mu^{2}-t})e^{-b\sqrt{n^{2}\mu^{2}-t}}.$ (15)
@@
The analysis of the high energy data on the proton-antiproton scattering in the framework of this model shows the obvious non-exponential behavior of the differential cross sections. For $\sqrt{s}=540$ GeV this shows the change in the slope from $16.8$ GeV-2 at $t=-0.001$ GeV2 up to $14.9$ GeV-2 at $t=-0.12$ GeV2. For the same $t$ the size of $\rho(s,t)$ changes from $0.141$ up to $0.089$. For Tevatron energy $\sqrt{s}=1800$ GeV it shows the change of the slope from $18.1$ GeV-2 at $t=-0.001$ GeV2 up to $15.9$ GeV-2 at $t=-0.12$ GeV2 and again the size of $\rho(s,t)$ changes from $0.182$ up to $0.143$. Hence the model shows the continuously decreasing slope and $\rho$ at small $t$. Practically the same results were obtained in [36] in the framework of the model that also uses the eikonal unitarization.

### v3

@@ Abstract
In the framework of the Regge-eikonal model of hadron interaction based on the analyticity of the scattering amplitude with taking into account the hadron structure, a simultaneous analysis is carried out of 90 sets of data. These sets include the data were obtained at low energies ($\sqrt{s}>3.6$ GeV and high energies at FNAL, ISR, $SP\bar{P}S$, TEVATRON and LHC with 4326 experimental points, including the polarization data of analysing power. The energy and momentum transfer dependence of separate sets of data is analyzed on the basis of the eikonalized Born amplitude with taking into account two additional anomalous terms. Different origins of the nonlinear behavior of the slope of the scattering amplitude are compared. No contribution of hard-Pomeron in the elastic hadron scattering is found. The importance of Odderon’s contribution is presented. The form and energy dependence of different terms of the cross-even and cross-odd parts of the elastic nucleon-nucleon scattering amplitude is determined in the framework of the High Energy Generalize Structure (HEGS) model. In the framework of the HEGS model, using the electromagnetic and gravitomagnetic form factors, the differential cross sections in the Coulomb Nuclear Interference (CNI) region and at large momentum transfer are described well in a wide energy region simultaneously. It is shown that the cross-even part includes the soft pomeron and the additional term with a large slope have energy dependence $\ln^{2}(s)$. The cross-odd part includes the maximal odderon term with $\ln^{2}(s)$ and an additional oscillation term with $\ln(s)$. It is shown that the additional terms with large slope are proportional to charge distributions but the maximal odderon term and oscillation term are proportional to matter distributions. Also, a good description of proton-neutron differential scattering with 526 experimental points is obtained on the basis of the amplitudes taken from the analysis of $pp$ and $p\bar{p}$ scattering. A good enough description of the polarization data was also obtained.
@@ I Introduction
One of the most important tasks of modern physics is the research into the basic properties of hadron interactions. The dynamics of strong interactions finds its most complete representation in elastic scattering. It is just this process that allows the verification of the results obtained from the main principles of quantum field theory: the concept of the scattering amplitude as a unified analytic function of its kinematic variables connecting different reaction channels was introduced in the dispersion theory by N.N. Bogoliubov[1]. Now many questions of hadron interactions are connected with modern problems of astrophysics such as unitarity and the optical theorem [2], and problems of baryon-antibaryon symmetry and CP-invariance violation [3]. Many models predict that soft hadron interactions will enter a new regime at the LHC: given the huge energy, the S-matrix reaches the unitarity limit. The main domain of elastic scattering is small angles. Only in this region of interactions we can measure the basic properties that define the hadron structure. Their values are connected, on the one hand, with the large-scale structure of hadrons and, on the other hand, with the first principles that lead to theorems on the behavior of scattering amplitudes at asymptotic energies [4, 5].
@@
The research of the structure of the elastic hadron scattering amplitude at superhigh energies and small momentum transfer - $t$ can give a connection between the experimental knowledge and the basic asymptotic theorems based on first principles [6, 7, 5]. This gives information about the hadron interaction at large distances where the perturbative QCD does not work [8, 9, 11] and a new theory as, for example, instanton or string theories must be developed.
@@
The structure of hadrons reflected in generalized parton distributions (GPDs) is now one of the most interesting questions of the physics of strong interactions (see, for example, [13, 12]. This is tightly connected with the spin physics of the hadron [14]. Some modern accelerator experiments have developed extensive programs for deep studies of different issues related to this problem. For example, the Jefferson Laboratory/Electron-Ion Collider (EIC) team intends to extract generalized parton distributions (GPDs) [15]. The same problem is posed for future experiments at the SPD of NICA-JINR [16].
@@
Recent studies of elastic scattering of high energy protons lead to several unexpected results reviewed, e.g., in [9, 10]. Spin amplitudes of the elastic $NN$ scattering constitute a spin picture of the nucleon. Without knowing of the spin $NN$-amplitudes, it is impossible to understand the spin observable of nucleon scattering off nuclei. In the modern picture, the structure of hadrons is determined by GPDs which include the corresponding parton distributions (PDFs). The sum rules [17, 18, 19, 20] allow one to obtain the elastic form factor (electromagnetic and gravitomagnetic) through the first and second integration moments of GPDs [21]. This leads to remarkable properties of GPDs, some corresponding to inelastic and elastic scattering of hadrons. Now different models examining the nonperturbative instanton contribution lead to sufficiently large spin effects at superhigh energies [22, 23, 24, 25]. The research of such spin effects will be a crucial stone for different models and will help us to understand the interaction and structure of particles, especially at large distances. There are large programs of researching spin effects at different accelerators. Especially, we should like to note the programs at the NICA where the polarization of both the collider beams will be constructed. So it is very important to obtain reliable predictions for spin asymmetries at these energies. In this paper, we extend the model predictions to spin asymmetries in the NICA energy domain.
@@
There are two sets of data - at a small momentum transfer [26] and at a middle and large momentum transfer [27]. They overlap in some region of the momentum transfer, which supplies practically the same normalization of both sets of differential cross sections of elastic proton-proton scattering. Recently, the first set of data has created a wide discussion of the determination of the total cross section and the value of $\rho(t=0)$ [28].
@@
There is a very important characteristic of the elastic scattering amplitude such as the ratio of the real part to imaginary part of the scattering amplitude - $\rho(s,t)$. It is tightly connected with the integral and differential dispersion relations. Of course, especially after different results obtained by the UA4 and UA4/2 Collaborations, physicists understand that $\rho(s,t=0)$ is not a simple experimental value but heavily depends on theoretical assumptions about the momentum dependence of the elastic scattering amplitude. Our analysis of both experimental data obtained by the UA4 and UA4/2 Collaborations shows a small difference value of $\rho(s,t=0)$ obtained in both the experiments if the nonlinear behaviour of the elastic scattering amplitude is taken into account [29]. Hence, this is not an experimental problem but a theoretical one [30].
@@
For extraction of the sizes of $\sigma_{tot}$ and $\rho(t=0)$ the Coulomb hadron region of momentum transfer is used (for example [26]). However, the form of the scattering amplitude assumed for small $t$ and satisfying the existing experimental data at small momentum transfer can essentially be different from experimental data at large $t$. One should take into account the analysis of the differential cross section at 13 TeV where the diffraction minimum impacts the form of $d\sigma/dt$ already at $t=-0.45$ GeV2.
The analysis of new effects discovered on the basis of the experimental data at 13 TeV [30, 31, 32] and associated with the specific properties of the hadron potential at large distances was carried out with taking account all sets of experimental data on elastic $pp$-scattering obtained by the TOTEM and ATLAS Collaborations in a wide momentum transfer region and gave a quantitative description of all examined experimental data with minimum fitting parameters.
@@
The non-linear behavior of the slope of the differential cross sections at a small momentum transfer, which was announced by the TOTEM Collaboration in the proton-proton elastic scattering at $8$ TeV, shows that the complex (complicated) form of strong interactions searched out at low energies remains at superhigh energies too. This means that the strong interaction is not simplified at superhigh energies but includes many different parts of the hadron potential.
@@
Using the existing model of nucleon elastic scattering at high energies $\sqrt{s}=[3.6-14$] TeV [33, 34], which involves minimum of free parameters, we are going to develop its extended version aimed to describe all available data on cross sections and spin-correlation parameters at lower energies down to the SPD NICA region. The model will be based on the usage of known information on GPDs in the nucleon, electro-magnetic and gravitomagnetic form factors of the nucleon taking into account analyticity and unitarity requirements and providing compatibility with the high energy limit, where the pomeron exchange dominates.
The structure of the paper is as follows: in Section 2 the basic picture of diffraction scattering is discussed including the nonlinear behavior of the slope of the differential cross section at small momentum transfer. In Section 3 the experimental data base used in the analysis is reviewed. In section 4 the asymptotic structure of the scattering amplitude is discussed shortly and the form factors used are particularly considered. Section 5 presents the different model approximations used in the model and the structure of the scattering amplitude. In section 6 the results of the fitting procedure and the behavior of the imaginary and real parts of the elastic scattering amplitude are considered. In section 7 the impact of the value of $\rho(s,t)$ and odderon contributions on the differential cross sections is analysed. Section 8 discusses new effects discovered during the model analysis of experimental data. In section 9 the obtained results for proton-neutron elastic scattering are presented. In section 10 the results of the model calculations of the spin correlation parameter $A_{N}(s,t)$ are present. Finally, the obtained results and some predictions of our model are discussed in Section 11.
@@ II.1 Electromagnetic scattering
The $t$ dependence of the elastic peak is one of the important issues, which can also be related to optics. At high energy, elastic scattering at small angles has some similarity with light scattering in the Fresnel region. The amplitude for the electric field propagating in the $z$ direction is given by the Rayleigh-Sommerfeld equation
@@
where $r=\sqrt{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}+z^{2}}$,and $\cos\theta=\frac{z}{r}$. The integral can be performed analytically only for the simplest geometrical cases. At $x=kR>>1$ it can be obtained [35]
${\cal A}(\theta)\propto\frac{iJ_{1}(x\sin\theta)}{\sin\theta}$
Hence, even in the simplest case the elastic peak cannot be described by a simple exponential. Note that this form of the scattering amplitude is used in diffractive hadron scattering.
@@ II.2 Hadron scattering
In ref. [36], the TOTEM Collaboration announced the observation of the non-exponential behavior of the differential elastic cross sections at $8$ TeV and small momentum transfer $|t|$. Of course, the form of the elastic peak depends also on the structure of particles and the dynamics of the interaction. These two features can be parameterised by the profile function $\rho(\vec{b})$ (in the space of impact parameter, $\vec{b}$) or by the elastic form factor (in momentum space). For example, in [37] different forms of the profile function determined by the density distributions taking part in the interaction were considered:
$circle\ of\ radius\ b_{0}=4\ GeV^{-1}$
@@
$hollow\ disk\ near\ b_{0}=2.8\ GeV^{-1}$
@@
[…]
@@
Figure 1: Possible $t$ dependences of the scattering amplitude ${\cal A}(t)$, rescaled to unity at $t=0$. (solid line - for Eq. (5); long-dashed line - for Eq. (2); short-dashed line - for Eq. (4); dotted line - for Eq. (3), and dashed-dotted line - for Eq. (6).
The $t$ dependence of these amplitudes is shown in Fig. 1. Clearly, the simple exponential form (5) is a special separate case. We can see, except for the last case (eq.6), that if we take a small interval of $t$ (for example, $0.150.01<|t|<$ 0.15 GeV2) the curves give practically the same result, especially the lines of eqs. 3,5,6. However, they give essentially different results on some larger values of $t$ or as $t\rightarrow 0$. Hence, to obtain the true form of the scattering amplitude, it is necessary to take experimental data in a wider interval of $t$ and tending to the limit $t\rightarrow 0$.
This simplest approximation of the differential cross section by one exponent with a constant slope can lead to artificial new effects. For example, in the experiment at the Protvino accelerator on elastic proton-proton scattering there were found ”oscillations” at a small momentum transfer [38]. However, in [39] it was shown that such ”oscillations” can appear when the differential cross section is described by two exponentials with different slopes and one exponential with the constant slope.
@@ II.3 Non-linear slope
The complicated $t$ dependence of the slope can have many origins. First of all, it comes from the unitarization procedure of the Born term of the elastic scattering amplitude. In many purely phenomenological analyses it is represented by the $Ct^{2}$ term in the slope that mimics the unitarization procedure. The hadron structure is reflected also not only at a large momentum transfer but also at small $t$. This especially concerns the meson clouds whose interaction in many models is added to the central part of hadrons (for example, in the Pumplin model [40] and the Dubna Dynamical (DD) model [41] which leads to the $\sqrt{t_{0}+t}$ dependence of the scattering amplitude). Such a complicated hadron structure can be reflected in the presence of two form factors - electromagnetic and mater form factor of hadrons (for example, the HEGS model [33, 34]).
@@
is used in many phenomenological descriptions of the elastic differential cross sections to explain the ”break” in the differential cross sections. Note that an additional term in the slope like $\sqrt{t_{0}-t}$ was obtained earlier whose first approximation can be related to absorbtion corrections that can produce a set of canceling Regge cuts [42, 43] and leads to the Schwarz type trajectories [44] $\alpha(t)=1+\gamma t^{1/2}$. A more complicated form was obtained in [45] $\alpha_{\pm}(t)=1\pm\gamma t^{1/2}+2\rho(1/2\ \gamma^{2}t)^{3/2}(-ln(t))^{1/2}$. The appearance of a complex trajectory greatly complicates the picture and requires additional research. Hence, in [46], based on the works [47, 48], it was proposed to use the simplest form $\alpha(t)=1.041-0.15\sqrt{t_{0}-t}$. However, as we show, such behavior has a really pure phenomenological basis and can be replaced by either a simpler form (for example, eq.(7) or a more complicated form used in the HEGS-model [34]).
Many models based on the famous works [49, 50] researched the non-linear behavior of the scattering amplitude. Based on the works in [51], it was obtained that
@@
with $q^{2}=-t$ Note that we have removed the misprint in this equation, as made in [52, 53]. They obtained the limits of the representation in the brackets at $q^{2}>>4\mu^{2}$; hence, the slope grows in order $q^{2}$ with small logarithmic suppression. At small $t$ ($q^{2}\ll 4\mu^{2}$ ) they predicted that the representation in the brackets goes to $ln(m^{2}/\mu^{2})-8/3$. Note that the authors aimed to explain the deviation of the slope from the constant at a non-small momentum transfer (in the domain $-t=0.4$ GeV2). However, in this domain the impact of the diffraction minimum is already felt. Hence this domain of $t$ is usually described by an additional term like $ct^{2}$, which were proposed earlier (for example [54]).
Figure 2: The $t$ dependence of the different parts of the terms of eq.(9) (solid line - the sum of all three terms in the brackets; the dots-dashed line - the first term is multiplied by ($-1$); the long-dashed line - the second term, the short-dashed line - the sum of the two first terms).
Figure 3: The $t$ dependence of the additional terms to $\alpha^{\prime}t$ (solid line - eq.(9); long-dashed line - eq.(7); short-dashed line - eq.(7) without extraction of $2\mu$ and multiplied by $0.1$.
@@
In [52], using the calculations [51] and removing the misprint, the authors proposed for $h_{1}$, with taking into account the meson form factor (with $\Lambda_{\pi}^{2}=m_{\rho}^{2}$ GeV2 ), the following equation:
$h_{1}(q^{2})=$ $\frac{4}{\tau}f^{2}_{\pi}(t)[2\tau-(1+\tau)^{3/2}$ (12)
$ln\frac{\sqrt{1+\tau}+1}{\sqrt{1+\tau}-1}$ $+ln\frac{m^{2}}{\mu^{2}}],$
@@
Really, for the slope, which is multiplied by $q^{2}$, there are two divergence terms with different signs plus a small constant term. The divergence terms cancel each other and the rest have a slow $t$ dependence (see Fig. 2). The cancelation of the two terms as $t\rightarrow 0$ is not full and the rest have also some indefiniteness and strongly depend on $t$. In Fig. 2, the $t$ dependence of the different parts of the term of eq.(9) in the brackets is shown. The cancelation of the two diverse terms leads to the negative contribution to the standard constant slope. However, the sum of all three terms gives a positive additional contribution with a small $t$ dependence.
@@
In Fig. 3, we show the $t$ dependence of a different form of the slopes with the full kinematic coefficient. One can see a very different $t$ dependence , especially in the form $\sqrt{4\mu^{2}+q^{2}}$ which was proposed in [46] without the extraction of $2\mu$. Some models of hadron interactions at high energy suppose that the slope has a slow $t$ dependence. For example, the Dubna dynamical model (DDm) [41], which takes into account the contribution to hadron interactions from the meson cloud of the nucleon and uses the standard eikonal form of the unitarization, leads to the scattering amplitude in the form
$T(s,t)=$ $-is\sum_{n=1}^{\infty}\frac{\mu}{(n^{2}\mu^{2}-t)^{3/2}}$
$(1-b\sqrt{n^{2}\mu^{2}-t})e^{-b\sqrt{n^{2}\mu^{2}-t}}.$
@@
The analysis of the high energy data on the proton-antiproton scattering in the framework of this model shows the obvious non-exponential behavior of the differential cross sections. For $\sqrt{s}=540$ GeV it shows the change in the slope from $16.8$ GeV-2 at $t=-0.001$ GeV2 up to $14.9$ GeV-2 at $t=-0.12$ GeV2. For the same $t$ the size of $\rho(s,t)$ changes from $0.141$ up to $0.089$. For Tevatron energy $\sqrt{s}=1800$ GeV it shows the change of the slope from $18.1$ GeV-2 at $t=-0.001$ GeV2 up to $15.9$ GeV-2 at $t=-0.12$ GeV2 and again the size of $\rho(s,t)$ changes from $0.182$ up to $0.143$. Hence, the model shows the continuously decreasing slope and $\rho$ at small $t$. Practically the same results were obtained in [40] in the framework of the model that also uses the eikonal unitarization.